OpenStudy (anonymous):

(x + 8)3/2 = 343 A) {-1} B) {57} C) {41} D)

OpenStudy (anonymous):

d. \[\sqrt[3]{7}-8\]

OpenStudy (anonymous):

is it \[\frac{ (x+8) *3}{ 2 } = 343 ?\]

OpenStudy (anonymous):

no its a fraction 3/2

OpenStudy (anonymous):

ok same thing

OpenStudy (anonymous):